OpenStudy (anonymous):

average rate of change of f(x) from 4 to 9 f(x)x^2-7x

OpenStudy (anonymous):

Average rate of change \(\overline{f'}\) is \[ \overline{f'} = \frac{f(x_2)-f(x_1)}{x_2-x_1} \]

OpenStudy (anonymous):

In this case \((x_1,x_2)=(4,9)\)

OpenStudy (anonymous):

f(4)= (4)^2-7(4)=-12 f(9)=(9)^2-7(9)=18